# SPHERE: Spherical partitioning for large-scale routing optimization

### v1

@@ Abstract
We study shortest-path routing in large weighted, undirected graphs, where expanding search frontiers raise time and memory costs for exact solvers. We propose SPHERE, a source-target-aware heuristic that identifies an $s$-$t$ overlap: vertices that are close to both $s$ and $t$ in hop count. Selecting an anchor $a$ in this overlap partitions the task into two subproblems with unchanged problem-topology , $s\to a$ and $a\to t$; if either remains large, the procedure recurses on its induced subgraph. Because the cut lies inside the overlap, concatenating the resulting subpaths yields a valid $s\to t$ route without boundary repair. SPHERE is independent of the downstream solver (e.g., Dijkstra) and exposes parallelism across subproblems. On large networks, it achieves faster runtimes and smaller optimality gaps than Louvain-based routing and a METIS-based pipeline, even on graphs with more than a million nodes and edges, while also outperforming Dijkstra in runtime.
@@ 1 Introduction
Point-to-point routing in large weighted, undirected graphs is a canonical problem in operations research, one that emerged alongside the field’s inception.. Exact label-setting or label-correcting methods (e.g., Dijkstra) can certify optimality (Dijkstra, 1959), but their search frontiers grow quickly on large networks and become memory- and time-intensive. Graph partitioning methods (e.g., Kernighan-Lin, multilevel partitioning/METIS, spectral cuts) reduce problem size by cutting the graph into parts (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000), yet cuts chosen without regard to a specific query $(s,t)$ can intersect narrow passages that many feasible sstt routes must use, which can exclude high-quality routes or require costly boundary repair. Bidirectional search reduces depth by growing frontiers from $s$ and $t$ (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016), but still explores a single coupled state space and may expand many partial routes on large graphs. For completeness, we note that while feasibility for unconstrained shortest paths is trivial on connected graphs, constrained variants (e.g., resource- or budget-limited paths) are NP-hard; our focus here is the unconstrained case with nonnegative weights, which admits an optimal solution via Dijkstra’s algorithm (Dijkstra, 1959).
We propose SPHERE, a query-aware partition that aligns computation to the sstt geometry. Let $B_{s}(r)$ and $B_{t}(r)$ denote hop-distance spheres around $s$ and $t$. SPHERE starts from a radius tuple $(r_{s},r_{t})$ with guaranteed overlap $B_{s}(r_{s})\cap B_{t}(r_{t})\neq\emptyset$, then applies a monotone decrement rule that reduces the radii (when needed, both simultanously) so that $|r_{s}-r_{t}|\leq 1$ while preserving overlap, and selects the minimal balanced pair, i.e. no further admissible unit decrement of either radius (still keeping $|r_{s}-r_{t}|\leq 1$) leaves the intersection nonempty. Then it selects an anchor $a\in B_{s}(r_{s})\cap B_{t}(r_{t})$ and splits the query into $s\!\to\!a$ and $a\!\to\!t$, solved on the corresponding induced subgraphs under a tunable size cap. If a subproblem still exceeds the cap, the procedure recurses on its induced subgraph. Because the cut lies inside the overlap, concatenating subpaths at $a$ yields a valid sstt route without boundary repair. The partition is solver-agnostic and the independent subproblems can be solved in parallel.
Our contributions are:
@@
    Query-aware partition. We introduce a simple heuristic that adaptively splits the graph based on the geometry of the source-target pair, ensuring that each part preserves connectivity and can be rejoined without additional repair.
@@
    Parallel subinstances. The subproblems generated by SPHERE are independent induced subgraphs, naturally exposing parallelism and enabling scalable routing on distributed systems with effort scaling to hardware.
@@
    Empirical validation. On large road graphs, SPHERE achieves faster runtimes than Louvain-based routing and a METIS-based pipeline while matching or improving route quality, benefiting further from inherent parallelism.
SPHERE is complementary to downstream solvers and can wrap Dijkstra or A* (Dijkstra, 1959; Hart et al., 1968), as well as modern speedup indices such as ALT/landmarks (Goldberg and Harrelson, 2005), reach-based pruning (Gutman, 2004), contraction hierarchies (Geisberger et al., 2008), hub labeling (Abraham et al., 2011), and transit-node routing (Bast et al., 2016); we evaluate with common solvers and baselines and discuss integration in Section 5. Next, we motivate these design choices by examining where existing approaches trade-off feasibility, and instance size, before positioning SPHERE within the literature and formalizing our approach.
@@ 2 Motivation
Routing at scale must meet three criteria simultaneously: control per-instance complexity, make an $s$-$t$ route feasible, and remain compatible with trusted downstream solvers while enabling parallel execution. Existing tools typically sacrifice at least one of these: Labeling methods (Feillet et al., 2004; Irnich and Desaulniers, 2005) preserve feasibility but generate a combinatorial explosion of partial routes on large graphs. Graph partitioners (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000) control instance size, but because their boundaries are chosen without regard to $s$ and $t$, they may cut across a bottleneck that every feasible sstt path must traverse, risking the loss of valid routes. Bidirectional search (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016) reduces search depth by expanding from both ends, $s$ and $t$, but the two frontiers still form a single combined search space and therefore cannot be partitioned into smaller, independent pieces suitable for parallel solution.
SPHERE overcomes these trade-offs by aligning the partition with the specific request. To ensure controllable size, it expands hop-distance spheres from $s$ and $t$ until reaching the last nonempty overlap, where a user-defined radius limits the diameter of each subproblem. Second, for feasibility, SPHERE selects an anchor $a$ in the overlap and solves $s\!\to\!a$ and $a\!\to\!t$ on their respective induced subgraphs, so cuts always lie inside the overlap and concatenation at $a$ is guaranteed to generate a feasible sts-t route. This preserves all local connectivity within each part and ensures that every feasible subpath remains feasible in the global graph. Third, the resulting subproblems are independent, enabling parallel execution across multiple cores, and the design is solver-agnostic which means that any downstream optimizer, from exact Mixed-Integer Programming (MIP) on small pieces to strong metaheuristics on larger ones (Pisinger and Ropke, 2007; Vidal, 2022), can be applied without modification. In short, SPHERE offers a partition strategy that controls complexity, preserves feasibility, and exposes natural parallelism while remaining compatible with the solvers practitioners already trust.
@@ 3 Related Work
Point-to-point routing on large weighted graphs has been studied from several perspectives. Classical exact methods such as label-setting and label-correcting algorithms (e.g., Dijkstra) can guarantee global optimality on connected graphs with nonnegative weights (Dijkstra, 1959). Dynamic-programming and labeling variants extend this to richer settings and can certify optimality, but they often face an explosion in the number of labels on large graphs (Feillet et al., 2004; Irnich and Desaulniers, 2005). Approximation schemes exist for restricted variants (Hassin, 1992), trading accuracy for speed, yet these techniques operate on the full graph and therefore do not reduce the instance size before optimization.
Graph partitioning offers a complementary idea: simplify the problem by splitting the graph. Kernighan-Lin improves an initial cut by greedy node swaps to reduce the number of edges crossing between parts while keeping parts balanced (Kernighan and Lin, 1970). Multilevel schemes such as METIS repeatedly coarsen the graph, compute a cut on the coarse instance, and refine it during uncoarsening, which gives strong scalability in practice (Karypis and Kumar, 1998). In routing pipelines, these partitions are often used to build a quotient graph for a coarse path that is later refined within parts. A limitation of all such methods is that the cut is chosen without regard to the specific start-target pair $(s,t)$, so the boundary can lie across a narrow passage that every feasible $s$-$t$ path must use. Once this happens, the coarse or restricted routing stage may miss or degrade the necessary connection.
Bidirectional search reduces search depth by expanding frontiers from both ends and meeting near the middle (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997). The MM algorithm of Holte et al. (2016) provides a priority rule that ensures expansions do not go beyond the midpoint of the optimal solution. These methods improve exploration efficiency but still operate within a single coupled search space and do not produce small, independent instances. In large graphs this can still generate many partial routes, and unlike partition-based methods, they do not offer a straightforward path to parallel, solver-agnostic execution.
@@ 4.1 Spherical subgraphs
We begin by fixing distance and the local “window” around a node. In the following, let every graph $G=(V,E)$ be a finite, connected, undirected graph with positive edge weights. Unless stated otherwise we measure distance in hops (unweighted shortest path); all statements below remain valid for the weighted shortest-path distance.
Definition 1 (Spheres).
@@
$S_{R}(\tilde{v})\;:=\;\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}\subseteq V,$
@@
Definition 2 (Spherical subgraph).
@@
$E^{\mathrm{sp}}_{R}(\tilde{v})=\bigcup_{v\in S_{R}(\tilde{v})}\mathcal{E}\!\bigl(\mathrm{SP}(\tilde{v},v)\bigr),$
@@
Definition 3 (Induced spherical subgraph).
@@ 4.2 Partition cut
With these building blocks in place, we now align the local windows to a specific query $(s,t)$. For radii $(R_{s},R_{t})\in\mathbb{N}^{2}$, define the overlap
$O(R_{s},R_{t})\;:=\;S_{R_{s}}(s)\cap S_{R_{t}}(t).$
We control how radii evolve, how they are initialized, and how an anchor is chosen in the overlap.
Definition 4 (Decrement rule).
A decrement rule is a map $\mathrm{decru}:\mathbb{N}^{2}\to\mathbb{N}^{2}$ that is monotone in the lexicographic order, i.e.,
@@
Example (monotone decrement).
A simple rule that satisfies the definition is
@@
It is coordinatewise nonincreasing, and it is lex-monotone: if $(R_{s},R_{t})\leq_{\mathrm{lex}}(R^{\prime}_{s},R^{\prime}_{t})$, then either $R_{s}<R^{\prime}_{s}$ (so the first branch cannot produce a larger first coordinate) or $R_{s}=R^{\prime}_{s}$ with $R_{t}\leq R^{\prime}_{t}$, in which case both images either decrement the same coordinate or the smaller pair cannot “overtake” the larger one in lexicographic order.
The map $\mathrm{decru}$ from Definition 4 determines how radii shrink once chosen. To make the procedure operative, we also need: (i) an initial choice of radii that guarantees the spheres overlap at the start, and (ii) a way to convert any nonempty overlap into a concrete split point. These are captured by the following two rules.
Definition 5 (Starting rule).
A starting rule $\mathrm{staru}:V\times V\to\mathbb{N}^{2}$ returns $(R_{s}^{0},R_{t}^{0})$ with nonempty overlap $O(R_{s}^{0},R_{t}^{0})\neq\emptyset$ (e.g., $R_{s}^{0}{+}R_{t}^{0}\geq d_{G}(s,t)$; a stronger but safe choice is $S_{R_{s}^{0}}(s)\cup S_{R_{t}^{0}}(t)=V$).
Definition 6 (Anchor rule).
An anchor rule $\mathrm{anru}:\mathcal{P}(V)\to V$ selects $a\in O(R_{s},R_{t})$ (e.g., uniform random).
Bringing these pieces together, the three rules: starting rule for initialization (Definition 5), decrement rule for shrinking (Definition 4), and anchor rule for selecting the split point (Definition 6), jointly define the following algorithm, which we refer to as the Partition cut:
@@
The intuition behind Algorithm 1 can be explained as follows and is visualized in Figure 2. Grow spheres around $s$ and $t$ until they overlap, then shrink via $\mathrm{decru}$ until the next shrink would empty the overlap. Pick an anchor $a$ according to Definition 6 in this last nonempty overlap, and split the query into $s\!\to\!a$ and $a\!\to\!t$ subproblems on the corresponding induced spheres.
@@
Figure 2: Last nonempty overlap. Blue and red spheres overlap in the green region containing both $a$ and $a^{\prime}$. Potential anchor node $a$ is selected according to Definition 6.
A natural question is whether the Partition cut of Algorithm 1 actually preserves feasibility of the original $s{\to}t$ routing problem. In other words, does splicing the two subpaths at the chosen anchor still yield a valid global route? The following statement makes this precise:
Proposition 4.1 (Feasibility preservation).
@@
Proof sketch.
Both subgraphs are induced, hence every used edge is in $G$. Since $a\in O(\bar{R}_{s},\bar{R}_{t})$, concatenating at $a$ yields a valid $s\to t$ path in $G$. ∎

### v2

@@ Abstract
We study shortest-path routing in large weighted, undirected graphs, where expanding search frontiers raise time and memory costs for exact solvers. We propose SPHERE, a query-aware partitioning heuristic that adaptively splits the problem by identifying source-target (sstt) overlaps of hop-distance spheres. Selecting an anchor node $a$ within this overlap partitions the task into independent induced subgraphs for $s\to a$ and $a\to t$, each restricted to its own induced subgraph. If resulting subgraphs remain large, the procedure recurses on that specific subgraph. We provide a formal guarantee that by using the partition cut within the shared overlap, the resulting subpaths preserve feasibility, thereby avoiding the need for boundary repair. Furthermore, SPHERE acts as a solver-agnostic framework that naturally exposes parallelism across subproblems. On million-scale road networks, SPHERE achieves faster runtimes and smaller optimality gaps than contemporary state-of-the-art partitioning and community-based routing pipelines. Crucially, it also substantially mitigates heavy-tail runtime outliers suffered by standard exact methods, yielding highly stable and predictable execution times across varying queries.
@@ 1 Introduction
Point-to-point routing in large weighted, undirected graphs is a canonical problem in operations research, one that emerged alongside the field’s inception. Exact label-setting or label-correcting methods (e.g., Dijkstra) can certify optimality [Dijkstra1959], but their search frontiers grow quickly on large networks and become memory- and time-intensive. Graph partitioning methods (e.g., Kernighan-Lin, multilevel partitioning/METIS, spectral cuts) reduce problem size by cutting the graph into parts [KernighanLin1970, KarypisKumar1998, ShiMalik2000], yet cuts chosen without regard to a specific query $(s,t)$ can intersect narrow passages that many feasible sstt routes must use, which can exclude high-quality routes or require costly boundary repair. Bidirectional search reduces depth by growing frontiers from $s$ and $t$ [Nicholson1966, Pohl1969, KaindlKainz1997, Holte2016], but still explores a single coupled state space and may expand many partial routes on large graphs. For completeness, we note that while feasibility for unconstrained shortest paths is trivial on connected graphs, constrained variants (e.g., resource- or budget-limited paths) are NP-hard; our focus here is the unconstrained case with nonnegative weights, which admits an optimal solution via Dijkstra’s algorithm [Dijkstra1959].
We propose SPHERE, a query-aware source-target partitioning heuristic that aligns computation directly with the sstt geometry. Let $S_{R}(s)$ and $S_{R}(t)$ denote hop-distance spheres centered around $s$ and $t$. SPHERE initializes a radius tuple $(R_{s},R_{t})$ large enough to guarantee a nonempty overlap; $S_{R_{s}}(s)\cap S_{R_{t}}(t)\neq\emptyset$. SPHERE then applies a monotone decrement rule to systematically shrink the radii, maintaining a balanced scale ($|R_{s}-R_{t}|\leq 1$) until reaching the minimal balanced pair which refers to the situation where any further valid reduction of either radius would leave the intersection empty.
From this minimal overlap, the algorithm selects an anchor node $a\in S_{R_{s}}(s)\cap S_{R_{t}}(t)$. This anchor effectively splits the original source-target pair sstt into two independent subproblems, $s\to a$ and $a\to t$, which are then solved on their respective induced subgraphs subject to a tunable bound. This tunable bound is a bound that limits the maximum hop-radius of a subproblem which we refer to as $R_{\max}$. If a resulting subproblem still exceeds this cap, the procedure simply recurses on its induced subgraph.
Crucially, because the anchor $a$ sits inside the shared overlap, it exists in both induced subgraphs simultaneously. This guarantees that the local $s\to a$ path and the local $a\to t$ path will physically connect at $a$, seamlessly forming a continuous, valid sstt route without requiring any secondary algorithms to repair broken links across the cut. Finally, SPHERE is solver-agnostic and generates independent subproblems that can be executed in parallel. We make the following contributions:
@@
    Query-Aware Geometric Partitioning: We introduce a heuristic that adaptively partitions the graph based on the specific sts-t geometry by cutting inside the last nonempty overlap of hop-distance spheres. This ensures that each part preserves local connectivity and can be rejoined at an anchor node without the need for costly boundary repair.
@@
    Theoretical Guarantees: We formalize the spherical partitioning algorithm and prove that our cut placement intrinsically preserves global route feasibility. Furthermore, for unweighted, undirected graphs we establish theoretical conditions under which the recursive application of this partition guarantees a globally optimal shortest path.
@@
    Solver-Agnostic Parallel Execution: The subproblems generated by SPHERE are completely independent induced subgraphs. This allows the subproblems to be solved in parallel. Because SPHERE generates standard graphs, start nodes, and anchor nodes to the downstream solver, it is completely solver-agnostic. Any off-the-shelf routing algorithm (e.g. Dijkstra) can be used to compute local paths without the need to adjust the solver’s underlying structure.
  4. 4.
    Empirical validation. On large-scale networks with over a million nodes and edges, SPHERE achieves faster runtimes and smaller optimality gaps than other state-of-the-art sstt routing approaches. Crucially, it also substantially mitigates heavy-tail runtime slowdowns observed in standard methods, resulting in highly stable and predictable execution times across varying queries.
@@ 2 Motivation
Algorithms for routing at scale inherently struggle to balance search space complexity, route exactness, and practical scalability [bast2016route]. Building on these established challenges, routing at scale must meet three criteria simultaneously: control per-instance complexity, guarantee route feasibility, and enable parallel execution with standard solvers. Existing tools typically sacrifice at least one of these:
Labeling methods preserve feasibility, but generate a combinatorial explosion of partial routes due to the complexity of the topology of large graphs [Feillet2004, IrnichDesaulniers2005]. Graph partitioners such as [KernighanLin1970, KarypisKumar1998, ShiMalik2000] successfully control instance size, but their static boundaries ignore the specific $s$ and $t$ locations. This risks cutting across critical bottlenecks and losing feasible routes. Bidirectional search [Nicholson1966, Pohl1969, KaindlKainz1997, Holte2016] reduces depth but explores a single coupled state space that cannot be partitioned into independent pieces, making it difficult to enforce parallelism. While static partitioning methods like Corridor routing could theoretically be parallelized, they often require shared boundary information or synchronization across a global quotient graph to ensure path consistency. In contrast, SPHERE’s subproblems are mathematically independent by construction. Because the anchor node $a$ is guaranteed to exist in both induced subgraphs, no inter-process communication or boundary repair is required during the solve phase, allowing for true solver-agnostic parallel execution.
In comparison, SPHERE avoids combinatorial explosion by bounding subproblems within hop-distance spheres, strictly limited by a tunable bound ($R_{\max}$). Additionally, SPHERE cuts exclusively inside the shared sstt overlap and preserves local connectivity and guarantees a feasible route upon concatenation without any need for boundary repair. The induced subgraphs generated by SPHERE naturally expose parallel execution and act as a solver-agnostic framework, allowing practitioners to drop-in any trusted downstream optimizer, from exact Mixed-Integer Programming (MIP) to strong metaheuristics [PisingerRopke2007, Vidal2022], without modification.
In short, SPHERE offers a partition strategy that controls complexity, preserves feasibility, and exposes natural parallelism while remaining compatible with the solvers practitioners already trust.
@@ 3 Related Work
Point-to-point routing on large weighted graphs has been studied from several perspectives. Classical exact methods such as label-setting and label-correcting algorithms (e.g., Dijkstra) can guarantee global optimality on connected graphs with nonnegative weights [Dijkstra1959]. Dynamic-programming and labeling variants extend this to richer settings and can certify optimality, but they often face an explosion in the number of variables on large graphs [Feillet2004, IrnichDesaulniers2005]. Approximation schemes [Hassin1992] exist, trading accuracy for speed, yet these techniques operate on the full graph and therefore do not reduce the instance size before the optimization.
3.1 Graph Partitioning and Parallel Solvers
Graph partitioning offers a complementary idea: simplify the problem by splitting the graph, a strategy specifically suited for large-scale routing. Kernighan-Lin improves an initial cut by greedy node swaps to reduce the number of edges crossing between parts while keeping parts balanced [KernighanLin1970]. Multilevel schemes such as METIS repeatedly coarsen the graph, compute a cut on the coarse instance, and refine it during uncoarsening, which gives strong scalability in practice [KarypisKumar1998]. In routing pipelines, these partitions, along with community-based heuristics like the Louvain method, are often used to build a quotient graph for a coarse path or corridor that is later refined within parts [Blondel2008]. A limitation of all such methods is that the cut is chosen without regard to the specific start-target pair $(s,t)$, so the boundary can lie across a narrow passage that every feasible sstt path must use. Once this happens, the coarse or restricted routing stage may miss or degrade the necessary connection leading to potentially infeasible routes.
Parallel computation for shortest-path problems has also been studied to accelerate routing. For instance, Rozhon2023 analyze parallel Breadth-First Search and approximate distances within subgraphs. However, this approach is mainly theoretically analyzed.
3.2 Machine Learning and Deep Learning-Based Approaches
Recent advancements in neural combinatorial optimization have introduced Deep Learning models capable of solving complex routing tasks. Notable examples include Policy Optimization with Multiple Optima (POMO) [Kwon2020] and Multi-Task Vehicle Routing Solver with Mixture-of-Experts (MvMOE) [Zhou2024]. These neural solvers are primarily designed for distribution-trained variants of the Traveling Salesperson Problem (TSP) and Vehicle Routing Problem (VRP). Those methods operate effectively on small to medium-scale graphs.
3.3 Bidirectional Search
Bidirectional search reduces search depth by expanding frontiers from both ends and meeting near the middle [Nicholson1966, Pohl1969, KaindlKainz1997]. The MM algorithm of Holte2016 provides a priority rule that ensures expansions do not go beyond the midpoint of the optimal solution. These methods improve exploration efficiency but still operate within a single coupled search space and do not produce small, independent instances. In large graphs this can still generate many partial routes, and unlike partition-based methods, they do not offer a straightforward path to parallel, solver-agnostic execution.
@@ 4.1 Spherical subgraphs
We begin by formalizing the local "window" around a node using hop-based distance. Throughout this work, let $G=(V,E)$ denote a finite, connected, and undirected graph with positive edge weights. Unless otherwise specified, we measure distance in hops (unweighted shortest path), though our formulations remain valid for weighted shortest-path distances. For a comprehensive summary of the mathematical notation used in this paper, please refer to Appendix A.1.
Definition 4.1 (Spheres)
@@
$S_{R}(\tilde{v})\;\coloneqq\;\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}\subseteq V,$
@@
Definition 4.2 (Spherical subgraph)
@@
$E^{\mathrm{sp}}_{R}(\tilde{v})\coloneqq\Bigl\{\{i,j\}\in E:$ $d_{G}(\tilde{v},i)\leq R,\;d_{G}(\tilde{v},j)\leq R,$
$\bigl|d_{G}(\tilde{v},i)-d_{G}(\tilde{v},j)\bigr|=1\Bigr\}.$
@@
Definition 4.3 (Induced spherical subgraph)
[…]
@@ 4.2 Partition cut
With these geometric building blocks established, we now align these local windows to a specific query $(s,t)$.
Definition 4.4
For radii $(R_{s},R_{t})\in\mathbb{N}^{2}$, define the overlap as
$O(R_{s},R_{t})\;\coloneqq\;S_{R_{s}}(s)\cap S_{R_{t}}(t).$
In order to identify a precise splitting point, we control how the radii evolve from their initial state to a minimal configuration. This necessitates defining the properties of decrement rules, which systematically shrink the search space.
Definition 4.5 (Decrement rule)
A decrement rule
$\mathrm{decru}:\mathbb{N}^{2}\to\mathbb{N}^{2}$
is a map that is monotone in the lexicographic order, i.e.,
@@
Example 4.6 (monotone decrement)
A standard rule that satisfies these properties is one that prioritizes shrinking the larger sphere to keep the search balanced:
@@
This rule ensures the radii remain balanced ($|R_{s}-R_{t}|\leq 1$) throughout the process. If $(R_{s},R_{t})\leq_{\mathrm{lex}}(R^{\prime}_{s},R^{\prime}_{t})$, the smaller pair cannot "overtake" the larger one, preserving the order of operations.
We make this procedure operational by defining two auxiliary rules for initialization and anchor selection:
Definition 4.7 (Starting rule)
A starting rule
$\mathrm{staru}:V\times V\to\mathbb{N}^{2}$
returns $(R_{s}^{0},R_{t}^{0})$ with nonempty overlap $O(R_{s}^{0},R_{t}^{0})\neq\emptyset$ (e.g., $R_{s}^{0}{+}R_{t}^{0}\geq d_{G}(s,t)$; a stronger but safe choice is $S_{R_{s}^{0}}(s)\cup S_{R_{t}^{0}}(t)=V$).
Definition 4.8 (Anchor rule)
An anchor rule
$\mathrm{anru}:\mathcal{P}(V)\to V$
selects $a\in O(R_{s},R_{t})$ (e.g., uniform random).
The starting rule for initialization (Definition 4.7), decrement rule for shrinking (Definition 4.5) and anchor rule for selecting the split point (Definition 4.8), together define Algorithm 1, which we refer to as the Partition cut.
@@
The core intuition behind the Partition Cut (visualized in Figure 2) is to identify the smallest possible geometric region where the influence of $s$ and $t$ meet. Initially, spheres are grown around the terminals until they intersect. We then apply the decrement rule, $\mathrm{decru}$, to systematically shrink these radii until any further reduction would result in an empty intersection.
By selecting an anchor node $a$ within this final, minimal overlap (Definition 4.8), we effectively decompose the original $s\to t$ query into two independent $s\to a$ and $a\to t$ subproblems. These subproblems are then solved within their respective induced spherical subgraphs.
@@
Figure 2: Last nonempty overlap. Blue and red spheres overlap in the green region containing both $a$ and $a^{\prime}$. Potential anchor node $a$ is selected according to Definition 4.8.
A natural question is whether the Partition cut of Algorithm 1 actually preserves feasibility of the original $s{\to}t$ routing problem. In other words: Does splicing the two subpaths at the chosen anchor still yield a valid global route? The following statement makes this precise:
Proposition 4.9 (Feasibility preservation)
@@
Proof 4.10.
We provide a short proof sketch: Both subgraphs are induced, hence every used edge is in $G$. Since $a\in O(\bar{R}_{s},\bar{R}_{t})$, concatenating at $a$ yields a valid $s\to t$ path in $G$.